# Stationary states of forced two-phase turbulence

### v1

@@ 6.2 Parameters studied
For cases B1 and B2, the Weber number values are reduced to $9.75$ and $1.95$, respectively, by increasing $\sigma$ values to $0.158$ and $0.79$, from $0.079$ in case A. The Weber number was not further reduced because there was no breakup observed in case B2 [see Figure 12(c)]. For cases D1-D4, the density of phase 2 was varied from $0.01$ to $100$ while holding the density of phase 1 constant at a value of $1$ to achieve different density ratios. For cases E1-E4, the viscosity of phase 2 was varied while holding the viscosity of phase 1 constant to achieve different viscosity ratios. In addition, the density ratio was also varied in such a way that the kinematic viscosity is maintained the same in both phases and across cases E1-E4.
Case $\rho_{1}$ $\rho_{2}$ $\rho_{2}/\rho_{1}$ $\mu_{1}$ $\mu_{2}$ $\mu_{2}/\mu_{1}$ $\langle\phi_{2}\rangle$ $\langle\phi_{d}\rangle$ $\sigma$ $We_{L}$
A $1$ $1$ $1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.0655 0.0655 0.079 19.5
@@
C1 $1$ $1$ $1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.221 0.221 0.079 19.5
C2 $1$ $1$ $1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.524 0.476 0.079 19.5
C3 $1$ $1$ $1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.738 0.262 0.079 19.5
@@
D1 $1$ $10^{-2}$ $0.01$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.0655 0.0655 0.079 19.5
D2 $1$ $10^{-1}$ $0.1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.0655 0.0655 0.079 19.5
D3 $1$ $10^{1}$ $10$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.0655 0.0655 0.079 19.5
D4 $1$ $10^{2}$ $100$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-3}$ $1$ 0.0655 0.0655 0.079 19.5
@@
E1 $1$ $10^{-3}$ $0.001$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-6}$ $0.001$ 0.0655 0.0655 0.079 19.5
E2 $1$ $10^{-1}$ $0.1$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-4}$ $0.1$ 0.0655 0.0655 0.079 19.5
E3 $1$ $10^{1}$ $10$ $2.0334\times 10^{-3}$ $2.0334\times 10^{-2}$ $10$ 0.0655 0.0655 0.079 19.5
E4 $1$ $10^{3}$ $1000$ $2.0334\times 10^{-3}$ $2.0334$ $1000$ 0.0655 0.0655 0.079 19.5
@@ 7.1 Effect of Weber number
One of these two factors could be dominant, or both factors could play a role, depending on the state of the system, and be responsible for the increase in the specific kinetic energy of the dispersed phase with an increase in the Weber number.
@@ 7.1.2 Size distribution
where $We_{crit}=3$ is the critical Weber number, chosen based on Riviere et al. (2021). For both Weber numbers, the distribution follows a $-3/2^{th}$ power law for sub-Hinze scales and a $-10/3^{rd}$ power law for the super-Hinze scales, as seen before (Deane & Dale Stokes, 2002; Mukherjee et al., 2019; Riviere et al., 2021; Crialesi-Esposito et al., 2022; Roccon et al., 2023). In the stationary size distribution of the dispersed cloud, we can see that for $We_{L}=19.5$, the number of sub-Hinze scale bubbles is higher compared to the case with $We_{L}=9.75$ due to the increase in breakup at higher Weber numbers. This increase in the number of sub-Hinze scale bubbles will contribute to the higher specific kinetic energy of the dispersed phase in Figure 7(a) in terms of translational mode (Hypothesis 7.2), as described in Section 7.1.
@@ 8 Summary and conclusions
[…]
[Acknowledgements]This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. A preliminary version of this work has been published as a technical report in the annual publication of the Center for Turbulence Research (Jain & Elnahhas, 2023) and is available online [^1] https://web.stanford.edu/group/ctr/ResBriefs/2023/12_Jain.pdf.
[…]
[Funding]This work was partially supported by the startup grant to S. S. J. from the George W. Woodruff School of Mechanical Engineering at Georgia Institute of Technology. S. S. J. also acknowledge partial funding support from Boeing Co. and GWW-GTRI Connect seed grant. A. E. acknowledges support from NASA Transformation Tools and Technologies Grant #80NSSC20M0201.
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
[Author ORCIDs]
S. S. Jain, https://orcid.org/$0$-0002-9306-7903.
A. Elnahhas, https://orcid.org/$0$-0003-0449-9307.
[…]
[Author contributions]S.S.J. and A.E. conceptualized the study, S.S. J. performed the simulations, S.S.J. and A.E. performed the analysis, and wrote the original manuscript, including its revisions.

### v2

@@ 6.2 Parameters studied
For cases B1 and B2, the Weber number values are reduced to $9.75$ and $1.95$, respectively, by increasing $\sigma$ from case A. The Weber number was not further reduced because there was no breakup observed in case B2 [see Figure 12(c)]. For cases D1-D4, the density of phase 2 was varied while holding the density of phase 1 constant to achieve different density ratios. For cases E1-E4, the viscosity of phase 2 was varied while holding the viscosity of phase 1 constant to achieve different viscosity ratios. In addition, the density ratio was also varied in such a way that the kinematic viscosity is maintained the same in both phases and across cases E1-E4.
Case $\rho_{2}/\rho_{1}$ $\mu_{2}/\mu_{1}$ $\langle\phi_{2}\rangle$ $\langle\phi_{d}\rangle$ $We_{L}$
A $1$ $1$ 0.0655 0.0655 19.5
@@
C1 $1$ $1$ 0.221 0.221 19.5
C2 $1$ $1$ 0.524 0.476 19.5
C3 $1$ $1$ 0.738 0.262 19.5
@@
D1 $0.01$ $1$ 0.0655 0.0655 19.5
D2 $0.1$ $1$ 0.0655 0.0655 19.5
D3 $10$ $1$ 0.0655 0.0655 19.5
D4 $100$ $1$ 0.0655 0.0655 19.5
@@
E1 $0.001$ $0.001$ 0.0655 0.0655 19.5
E2 $0.1$ $0.1$ 0.0655 0.0655 19.5
E3 $10$ $10$ 0.0655 0.0655 19.5
E4 $1000$ $1000$ 0.0655 0.0655 19.5
@@ 7.1 Effect of Weber number
One of these two factors could be dominant, or both factors could play a role, depending on the state of the system, and be responsible for the increase in the specific kinetic energy of the dispersed phase with an increase in the Weber number. There is a small increase in the Taylor-scale Reynolds number of the carrier phase with $We_{L}$, which is due to the bubble/drop-induced agitation in the carrier phase.
@@ 7.1.2 Size distribution
where $We_{crit}=3$ is the critical Weber number, chosen based on Riviere et al. (2021). For both Weber numbers, the distribution follows a $-3/2^{th}$ power law for sub-Hinze scales and a $-10/3^{rd}$ power law for the super-Hinze scales, as seen before (Deane and Dale Stokes, 2002; Mukherjee et al., 2019; Riviere et al., 2021; Crialesi-Esposito et al., 2022; Roccon et al., 2023); although a more steeper power law for super-Hinze scales has also been observed in the past (Li et al., 2017) in more dynamic scenarios. In the stationary size distribution of the dispersed cloud, we can see that for $We_{L}=19.5$, the number of sub-Hinze scale bubbles is higher compared to the case with $We_{L}=9.75$ due to the increase in breakup at higher Weber numbers. This increase in the number of sub-Hinze scale bubbles will contribute to the higher specific kinetic energy of the dispersed phase in Figure 7(a) in terms of translational mode (Hypothesis 7.2), as described in Section 7.1.
@@ 8 Summary and conclusions
Acknowledgements
This work was partially supported by the startup grant to S. S. J. from the George W. Woodruff School of Mechanical Engineering at Georgia Institute of Technology. S. S. J. also acknowledge partial funding support from Boeing Co. and GWW-GTRI Connect seed grant. A. E. acknowledges support from NASA Transformation Tools and Technologies Grant #80NSSC20M0201. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. A preliminary version of this work has been published as a technical report in the annual publication of the Center for Turbulence Research (Jain and Elnahhas, 2023) and is available online [^1] https://web.stanford.edu/group/ctr/ResBriefs/2023/12_Jain.pdf.